What is the width of a single slit through which 520 nm green light passes to form a first diffraction minimum at an angle of 20.0∘?
Answer #1

Solution)

Given,

Wavelength, λ = 520 x 10^-9 m

For first diffraction minimum, we know

sin θ = λ / a

Width of slit, a=?

So, sin20=(520*10^-9 m)/a

Width, a= 1.52*10^-6 m

===========